OpenStudy (lncognlto):

A big bag of jellybeans has a total of 112 white, blue, and red jellybeans. There are 4 more than twice as many blue jellybeans than red jellybeans. There are 12 more white jellybeans than red jellybeans. How many blue, red, and white jellybeans are there in the bag? The phrasing of this question is killing me. I get W + B + R = 112 and W = R + 12, but I'm struggling with the other.

OpenStudy (lncognlto):

@iGreen.

OpenStudy (anonymous):

Yes.. W + B + R = 112 This is correct. R = 2B + 4 R = W + 12

OpenStudy (anonymous):

I believe that's correct..but I'm not sure how to go on from there..

OpenStudy (lncognlto):

Okay, thanks. I can work it from there by substituting those values back into the original equation.